OpenStudy (anonymous):

100ml of solution of HCl of 3.65 g L^-1 are mixed with 100Ml of NaOH containing 2 g L^-1 what is the pH of resulting Solution?

OpenStudy (anonymous):

@Kryten

OpenStudy (anonymous):

ok lets take it long way round gama - mass concentration gama = m / V m= gama*V m(HCl) = 3,65 g/dm3 * 100 * 10-3 dm3 m(HCl)= 0,365 g n= m/M n(HCl)= 0,365 g / 36,5 gmol-1 = 0,01 mol m(NaOH) = 2 g/dm3 * 100*10-3 dm3 m(NaOH) = 0,2 g n(NaOH) = 0,2 g / 40 gmol-1 = 0,005 mol and if n(HCl)/n(NaOH) = 1/1 = 1 from reaction HCl + NaOH -> NaCl + H2O than n(HCl) that remains is n(HCl) - n(NaOH) = 0,01 - 0,005 = 0,005 and n(HCl) = n (H+) c(H+) = 0,005 mol / 0,2 dm3 c(H+) = 0,025 M pH = - log [H+] pH = 1,6